(1) for the function f , defined by f(x) = 4x^3 - 4x^2 -7x -2 , there exist a point c €]-1/2 ,2[ satisfying f'(c) = 0
(2) for all even integral values of n, lim (x+1)^-n as x tends to infinity exists.
1
Expert's answer
2020-04-02T08:04:52-0400

1) "f'=12x^2-8x-7"

"f'=0\iff x_1=-\frac{1}{2}; x_2=\frac{7}{6}"

both points are in the interval ["-\frac{1}{2};2" ]

2) "\lim\limits_{x\to \infty}(x+1)^{-n}=\left\{\begin{matrix}
 0 & n>0 \\
 1 & n=0 \\
\infty& n<0
\end{matrix}\right."

the limit doesn't exist for all negative n